OpenStudy (anonymous):

can someone please help!!? Tina wants to save money for school. Tina invests $1,200 in an account that pays an interest rate of 6.25% How many years will it take for the account to reach $11,900? Round your answer to the nearest hundredth

OpenStudy (anonymous):

M = P( 1 + i )^n where M= Final Amount P= Principal i = interest and n = years

OpenStudy (anonymous):

okay.. so its like A=P(1+r)^t

OpenStudy (anonymous):

So fill in the amounts you know...

OpenStudy (welshfella):

Write I as a fraction